### v1

# Fine-Tuning Large Audio-Language Models with LoRA for Precise Temporal Localization of Prolonged Exposure Therapy Elements thanks: We thank National Science Foundation under Grant # 2326144

Abstract
Prolonged Exposure (PE) therapy is an effective treatment for post-traumatic stress disorder (PTSD), but evaluating therapist fidelity remains labor-intensive due to the need for manual review of session recordings. We present a method for the automatic temporal localization of key PE fidelity elements—identifying their start and stop times—directly from session audio and transcripts. Our approach fine-tunes a large pre-trained audio-language model, Qwen2-Audio, using Low-Rank Adaptation (LoRA) to process focused 30-second windows of audio-transcript input. Fidelity labels for three core protocol phases—therapist orientation (P1), imaginal exposure (P2), and post-imaginal processing (P3)—are generated via LLM-based prompting and verified by trained raters. The model is trained to predict normalized boundary offsets using soft supervision guided by task-specific prompts. On a dataset of 313 real PE sessions, our best configuration (LoRA rank 8, 30s windows) achieves a mean absolute error (MAE) of 5.3 seconds across tasks. We further analyze the effects of window size and LoRA rank, highlighting the importance of context granularity and model adaptation. This work introduces a scalable framework for fidelity tracking in PE therapy, with potential to support clinician training, supervision, and quality assurance.
Index Terms:
Prolonged Exposure therapy, audio-language models, LoRA, QLoRA, temporal localization, multimodal learning, PTSD treatment, therapy fidelity, soft supervision.
I INTRODUCTION
Post-Traumatic Stress Disorder (PTSD) is linked to multiple co-morbidities and can severely impact quality of life. Recovery is possible through treatments like Prolonged Exposure (PE) therapy, a gold-standard, evidence-based approach  [1]. However, trained PE clinicians are in short supply. Therapist fidelity—adherence to PE’s core components—is essential for effective outcomes [2], but assessing fidelity typically requires manual review of full sessions, making it time-consuming and hard to scale. While prior work has explored automated evaluation using speech and language processing [3, 4, 5, 6], most approaches focus on utterance-level classification rather than precise temporal localization of extended therapeutic phases—an essential requirement for PE fidelity assessment. Furthermore, the creation of large, finely annotated datasets remains a major bottleneck. This paper introduces a framework for second-level temporal localization of key PE therapy fidelity phases—specifically  [7, 8]: therapist orientation to imaginal exposure (P1), duration of imaginal exposure (P2), and post-imaginal processing (P3). Our approach leverages Qwen2 Audio [9], a model that jointly processes audio and transcript inputs, fine-tuned using Low-Rank Adaptation (LoRA) [10] and quantization techniques (QLoRA) [11] for computational efficiency. To address annotation challenges, we employ a soft-supervision pipeline in which initial timestamps are proposed by a large language model (LLM) using session transcripts and subsequently verified by trained raters. The model is trained on fixed-duration audio-transcript windows paired with task-specific prompts (e.g., “When did the Imaginal Exposure start?”) to predict the normalized offset of the target event. Accurate temporal resolution is therefore critical, as the PE manual specifies that the three major components of the session—the initial check-in (P1), the imaginal exposure (P2), and the processing of the imaginal exposure (P3)—should last approximately 10, 50, and 10 mins, respectively  [8, 7].
Our contributions are: (1) a method for fine-tuning Qwen2 Audio with LoRA for multimodal temporal localization in PE therapy; (2) a soft-supervision strategy that combines LLM-based annotation with trained rater verification; (3) an empirical analysis of how window duration and LoRA configurations affect localization accuracy; and (4) a demonstration of our method on real-world PE therapy data, achieving a mean absolute error (MAE) of 5.3s.
To our knowledge, this is among the first efforts to fine-tune audio-language models for precise temporal localization of psychotherapy fidelity metrics. Section II reviews related work, Section III describes the dataset and annotation process, Section IV details our methodology, Section V presents experimental setup, Section VI has the results and discussion, and Section VII concludes.
II RELATED WORK
Automated analysis of psychotherapy sessions aims to provide scalable feedback, support training, and enhance therapeutic outcomes. Our work leverages pre-trained audio-language models to temporally localize psychotherapy fidelity metrics in session audio using multimodal inputs, using soft-supervision, where LLM-proposed timestamps are clinically verified.
II-A Automated Evaluation of Psychotherapy
Prior work has automated psychotherapy evaluation using NLP and speech processing. For session-level assessment, Chen et al. [3] developed a CBT quality pipeline using linguistic features, Flemotomos et al. [12] created a platform for MI skill evaluation through audio-based MISC coding, and Lin et al. [6] mapped dialogue to Working Alliance metrics. Li et al. [4] classified client utterances in text counseling with BERT, while Lin et al. [5] applied topic modeling to understand session content. For specific behavior detection, researchers employed multimodal techniques with BERT, VGGish, and HuBERT for MI codes and empathy assessment [13, 14]; LSTMs with attention for addiction counseling behavior coding [15, 16, 17]; MEMMs for counselor reflection detection [18]; language synchrony for empathy [19]; behavior trajectories in couples therapy [20]; and LLM embeddings for depression outcome classification [21]. Prosodic analysis examined therapist speaking style [22], empathy modeling [23], and ChangeTalk valence prediction [24]. These studies primarily address utterance-level classification or segment-level scoring rather than precise temporal localization of broader fidelity metrics.
II-B Temporal Event Localization in Audio and Speech
Wu et al.’s FLAM [25] enables frame-wise open-vocabulary sound event detection through global and frame-level embeddings, while Mishra et al.’s SING [26] processes speech with Whisper and LLaMA to determine speaker identity and location. Both offer temporal localization capabilities but target different domains than psychotherapy fidelity metrics.
II-C Large Pre-trained Audio-Language Models
Recent audio-language models include Krishna D N’s [27] cross-modal attention between Wav2Vec2.0 and BERT for emotion recognition, FLAM [25] and SING’s [26] combination of speech understanding with LLMs, and Li et al.’s [28] CLAP models associating audio and text. Our approach uniquely applies these models to localize psychotherapy fidelity metrics through specialized annotation and prompting.
II-D Prompting and Few-Shot Learning in Multimodal Models
Prior multimodal prompting approaches include Tsimpoukelli et al.’s “Frozen” [29] using images as visual prefixes, Li et al.’s PT-Text [28] for “audio-free” CLAP tuning, text instruction prompts with LLMs in SING [26], LLM prompting for interpretation [6], and prompt engineering for data generation [30]. Our approach directly guides model attention during temporal fidelity metric localization.
II-E Annotation and Supervision Strategies
We adopt a soft-supervision pipeline where LLM-proposed timestamps are verified by experts, offering a middle ground between intensive manual annotation in psychotherapy research [3, 12, 31] and fully automated approaches like FLAM’s synthetic data generation [25] or CALLM’s LLM-based mock profile augmentation [30], balancing reduced expert burden with clinical validity.
II-F Research Contribution
While existing research has advanced psychotherapy analysis, multimodal learning, and temporal event detection, our work addresses temporal localization of psychotherapy fidelity metrics using pre-trained models, researcher-driven prompts, and soft-supervision—laying the foundation for fine-grained, scalable fidelity assessment in clinical contexts.
Figure 1: Timeline of a 70-minute PE therapy session showing typical segments for three fidelity metrics: P1 (orientation to imaginal exposure), P2 (imaginal exposure), and P3 (processing). Durations vary across sessions; this structure represents a common clinical pattern. All durations are in minutes.
III DATASET
We begin with a dataset comprises 318 PE therapy sessions (see Table II) recorded at Emory University. Each 60-90 minute session contains a therapist-client interaction. All data is handled within IRB-approved environments with no PII leaving secure storage. We annotate PE therapy protocol elements: orientation (P1), imaginal exposure (P2), and processing (P3). As shown in Figure 1, P1, P2, and P3 typically occur sequentially across a 60–90 minute PE session, motivating our timestamp prediction task. To minimize manual effort, we implemented a zero-shot LLM pipeline that processes transcripts to produce JSON-formatted protocol presence labels with timestamp boundaries. To validate these labels while ensuring scalability, 10% (30) of the sessions were reviewed by trained raters. These raters were specifically trained to identify the presence and temporal boundaries of the key therapeutic elements, a process which yielded 94.4% timestamp accuracy (within 5-10 seconds) and 97.7% label prediction accuracy. This approach demonstrates the feasibility of our framework without requiring costly expert clinician time for the verification step. Audio files were downsampled from 44.1-48 kHz to 16 kHz, normalized, and converted to .wav format. Transcripts were generated using Amazon HealthScribe with sentence-level timestamps and speaker labels. After filtering sessions exceeding 1.5 hours (limitation of transcription service) or having alignment issues, we retained 313 recordings split into 219 training, 46 validation, and 48 testing sessions, with balanced distribution of therapy phases across splits.
TABLE I: Example LLM-generated annotations with timestamps (in sec) and ratings for PE therapy phases.
ID Phase Description Start (s) Stop (s) Present
1 Therapist orients the client to the planned imaginal exposure. 4.17 37.27 Yes
2 Imaginal lasted about 30–45 minutes (or about 15 for final imaginal). 105.83 2123.92 Yes
3 Therapist processes the imaginal exposure with the client. 2123.92 3023.83 Yes
TABLE II: Summary statistics of the audio dataset. Durations are shown in HH:MM:SS format, with raw seconds in parentheses.
Statistic Value
Number of files 318
Total duration 342.11 hours
Min duration 00:19:30 (1170.75 s)
Max duration 01:31:12 (5472.38 s)
Average duration 01:04:32.9 (3872.89 s)
Std. deviation 00:15:38.5 (938.55 s)
Figure 2: Overview of our fidelity-aligned modeling pipeline. A task-specific prompt guides QLoRA fine-tuning on Qwen2 Audio using 30-120s audio-transcript windows randomly sampled around annotated start/stop points. The model is trained to predict normalized temporal offsets using clinically validated or synthetic labels.
TABLE III: Per-phase average, start, and end MAE (in seconds) for timestamp prediction on the test set. We run the experiments on three seeds (42, 78, 123) and report their Mean $\pm$ S.D.
Window Config P1 Avg P1 Start P1 End P2 Avg P2 Start P2 End P3 Avg P3 Start P3 End
30s Head Only $6.8\pm 0.1$ $6.4\pm 0.2$ $7.0\pm 0.3$ $7.2\pm 0.2$ $7.2\pm 0.2$ $7.3\pm 0.3$ $6.8\pm 0.2$ $7.7\pm 0.6$ $5.8\pm 0.4$
LoRA (r=2) $5.8\pm 1.7$ $6.0\pm 1.5$ $5.6\pm 2.2$ $5.1\pm 2.3$ $5.7\pm 2.7$ $4.4\pm 2.3$ $4.8\pm 2.0$ $5.2\pm 3.0$ $4.4\pm 1.9$
LoRA (r=4) $6.2\pm 1.3$ $6.2\pm 2.0$ $6.0\pm 1.2$ $5.4\pm 2.0$ $5.4\pm 2.2$ $5.5\pm 2.0$ $4.9\pm 1.5$ $5.2\pm 2.0$ $4.6\pm 1.4$
LoRA (r=8) $5.9\pm 1.4$ $5.5\pm 1.5$ $6.4\pm 1.4$ $5.0\pm 1.8$ $5.2\pm 1.8$ $4.9\pm 1.8$ $5.0\pm 0.5$ $5.5\pm 1.8$ $4.4\pm 0.8$
60s Head Only $12.2\pm 0.8$ $11.5\pm 1.0$ $12.9\pm 0.7$ $13.9\pm 0.5$ $13.4\pm 2.2$ $14.4\pm 1.2$ $13.7\pm 0.6$ $15.3\pm 1.1$ $12.2\pm 0.6$
LoRA (r=2) $11.3\pm 2.5$ $11.5\pm 2.9$ $11.2\pm 2.6$ $12.1\pm 3.1$ $12.8\pm 3.0$ $11.4\pm 3.3$ $12.1\pm 2.9$ $13.4\pm 2.6$ $10.8\pm 3.4$
LoRA (r=4) $11.9\pm 2.1$ $11.0\pm 2.1$ $12.7\pm 2.0$ $10.2\pm 2.0$ $10.5\pm 2.3$ $9.9\pm 2.1$ $11.6\pm 1.4$ $11.3\pm 1.0$ $11.8\pm 2.4$
LoRA (r=8) $9.9\pm 0.1$ $9.6\pm 0.8$ $10.2\pm 0.7$ $9.7\pm 0.6$ $10.2\pm 0.9$ $9.2\pm 0.4$ $10.0\pm 0.6$ $9.8\pm 1.6$ $10.1\pm 0.5$
120s Head Only $25.2\pm 2.2$ $25.0\pm 2.4$ $25.4\pm 2.0$ $27.4\pm 0.7$ $25.1\pm 1.3$ $29.6\pm 2.8$ $24.4\pm 2.7$ $28.1\pm 3.1$ $20.8\pm 2.4$
LoRA (r=2) $20.7\pm 2.1$ $20.8\pm 2.0$ $20.5\pm 2.5$ $18.0\pm 2.3$ $17.5\pm 2.6$ $18.5\pm 2.3$ $21.4\pm 1.2$ $22.6\pm 0.9$ $20.2\pm 1.8$
LoRA (r=4) $20.7\pm 1.8$ $21.4\pm 2.1$ $19.9\pm 2.1$ $18.8\pm 2.1$ $18.3\pm 3.7$ $19.2\pm 1.9$ $21.0\pm 1.2$ $20.3\pm 2.4$ $21.7\pm 0.5$
LoRA (r=8) $20.1\pm 1.4$ $20.3\pm 1.2$ $19.8\pm 1.8$ $20.7\pm 1.9$ $21.7\pm 2.3$ $19.6\pm 1.5$ $22.3\pm 1.3$ $23.5\pm 3.5$ $21.2\pm 1.7$
IV METHODOLOGY
IV-A Problem Formulation
[…]
The objective is to minimize the error between the predicted normalized offset $\hat{o}_{j}$ and the true normalized offset $o_{j}$.
IV-B Dataset and Soft-Supervision Labeling
Our processed dataset consists of 313 PE therapy sessions, preprocessed as described in Section III, and transcribed using Amazon HealthScribe, which provides sentence-level timestamps and speaker attribution. Ground-truth start and stop times for fidelity metrics P1, P2, and P3 were initially generated via soft supervision using Claude Sonnet 3.5, prompted to return timestamps in seconds as illustrated in Table I. These LLM-generated annotations were subsequently reviewed and corrected by trained trained raters to create verified ground-truth labels. All audio recordings were downsampled to 16kHz and split at the session level into train, validation, and test sets.
IV-C Model Architecture and Fine-tuning
We fine-tune Qwen2-Audio-7B-Instruct [9], a large audio-language model that processes interleaved audio and text inputs through an audio encoder and a language model backbone. For each trained rater-verified boundary time, we construct focused input windows of fixed durations (30, 60, or 120 sec) containing both audio and aligned transcript segments. To ensure model robustness and enable data augmentation, the true boundary is randomly positioned within each window by normalizing its location to the $[0,1]$ interval and adjusting the window center accordingly. The training target is the normalized offset of the boundary relative to the window start. Each input is prepended with a task-specific prompt, such as: “The following audio and transcript segment is focused around the START of ’Was Prolonged Exposure done in the session?’. Identify the normalized offset (between 0.0 and 1.0) of this precise START within the given segment.” We use QLoRA [11] for memory-efficient fine-tuning, applying 4-bit NormalFloat (NF4) quantization with bfloat16 computation, and LoRA [10] to the LLM component with ranks $r\in{2,4,8}$, scaling factor $\alpha=2r$, and dropout of 0.1. A regression head receives the final hidden state of the last non-padding token and predicts the normalized offset via a LayerNorm, two linear layers with ReLU activation, and a final Sigmoid function. All experiments were conducted using three random seeds (42, 78, and 123), and results are reported as mean $\pm$ standard deviation across these runs.
IV-D Training and Evaluation
The model minimizes MAE using the AdamW optimizer [32] (learning rate $1\times 10^{-4}$, weight decay 0.01, cosine schedule with 0.1 warmup ratio). Training runs for 10 epochs with a batch size of 1 and early stopping (patience = 3) based on validation MAE. We report MAE in seconds, calculated by denormalizing predicted offsets to absolute timestamps and comparing them to ground truth. Results are reported separately for each metric type (P1, P2, P3).
V EXPERIMENTAL SETUP
We implemented the temporal localization framework described in Section IV using Qwen2-Audio-7B-Instruct [9]. Experiments were conducted on a dataset of 313 PE therapy sessions, split into 219 train, 46 validation, and 48 test files as detailed in Section III. To assess the influence of context length and fine-tuning strategies, we varied (1) input audio-transcript window durations (30s, 60s, and 120s), and (2) LoRA ranks $r\in\{2,4,8\}$. These configurations were compared against a “Head Only” baseline where only the final regression layer was trained. All other QLoRA and training parameters (optimizer, learning rate, epochs, etc.) followed the specifications described in Section IV. Performance was evaluated using MAE in seconds for PE elements P1, P2, and P3.
VI RESULTS AND DISCUSSION
Table III presents the mean $\pm$ standard deviation of MAE across three random seeds. Our best-performing configuration—30-second windows with LoRA rank $r=8$—achieved an average MAE of 5.3 seconds compared to ground truth across P1 ($5.9\pm 1.4$s), P2 ($5.0\pm 1.8$s), and P3 ($5.0\pm 0.5$s). This level of precision is promising for downstream clinical use, where timestamp-level fidelity tracking is essential. Earlier attempts at feature-level fusion using independent audio and text encoders produced overlapping embeddings across therapeutic phases, limiting their temporal resolution. In contrast, the end-to-end audio-language architecture of Qwen2-Audio-7B enabled learning of joint contextual representations that more effectively captured therapeutic cues. Shorter input windows (30s) consistently resulted in lower MAEs compared to longer windows. For instance, with the “Head Only” model, P1 Avg MAE increased from $6.8\pm 0.1$s (30s) to $12.2\pm 0.8$s (60s) and $25.2\pm 2.2$s (120s). This trend held across all configurations and phases, suggesting that temporally focused input helps the model better localize fidelity events. LoRA fine-tuning generally outperformed the “Head Only” baseline, especially at 60s and 120s. At 30s, LoRA $r=2$ and $r=8$ both reduced MAE for P2 Avg to around 5.1s and 5.0s, respectively, compared to 7.2s with “Head Only”. At 60s, LoRA $r=8$ consistently delivered the best results across phases (e.g., P1 Avg: $9.9\pm 0.1$s, P2 Avg: $9.7\pm 0.6$s, P3 Avg: $10.0\pm 0.6$s) and showed strong stability with low standard deviations. For 120s windows, the trend was less consistent. LoRA $r=2$ yielded the lowest P2 Avg MAE ($18.0\pm 2.3$s), outperforming both $r=4$ and $r=8$. Similarly, for P3, LoRA $r=4$ slightly outperformed $r=8$. These results suggest that higher LoRA ranks may overfit when the input spans become too broad, and smaller adapters can generalize better in such settings. The optimal LoRA rank varied depending on both window size and phase, highlighting the need for task-specific tuning. LoRA models also tended to show higher variability across seeds compared to “Head Only”, though configurations like LoRA $r=8$ at 60s achieved both low MAE and strong consistency. Overall, these results reinforce the benefits of combining temporally focused inputs with appropriately tuned model adaptation for timestamp prediction in therapy analysis.
VII CONCLUSION
We introduced a QLoRA-based fine-tuning strategy for Qwen2-Audio to localize therapeutic elements in Prolonged Exposure sessions with high temporal precision. By leveraging both audio and transcript inputs within focused 30-second windows, our approach achieved an MAE of 5.3 seconds across rater-verified fidelity phases. Performance is strongly influenced by window size and LoRA rank, with shorter contexts and deeper adaptation yielding the most accurate predictions. These findings support scalable, automated fidelity assessment in PTSD therapy, aiding clinician supervision, training, and quality assurance. By localizing key therapeutic phases automatically, the approach reduces the burden of manual review. Future work will expand to additional fidelity markers and test generalization across diverse therapy settings.